HARDY CLASSES, INTEGRAL OPERATORS, AND DUALITY ON SPACES OF HOMOGENEOUS TYPE

STEVEN G. KRANTZ SONG-YING LI

Abstract. The authors study Hardy spaces, of arbitrary order, on a space of homogeneous type. This extends earlier work that treated only H^{p} for p near 1. Applications are given to the boundedness of certain singular integral operators, especially on domains in complex space.